OpenStudy (anonymous):

I need help please! Find the slope and the y-intercept of the equation 2x − 3y = 6.

OpenStudy (anonymous):

what are the answers

OpenStudy (anonymous):

A.)slope = 2/3 and y-intercept = −2 B.)slope = −2/3 and y-intercept = −2 C.)slope = 2/3 and y-intercept = 2 D.)slope = −2/3 and y-intercept = 2

OpenStudy (anonymous):

a